# TableBench: A Comprehensive and Complex Benchmark for Table Question Answering

### v1

@@ 4.3 Automatic Evaluation Metrics
we adopt Rouge-L (Lin, 2004) to assess the quality of the generated answers by measuring the n-gram overlap with reference answers. In the PoT method, we enforce a specific format for the executable code outputs and evaluate the final answer with the ROUGE-L metric, ensuring alignment with other reasoning methodologies. Specifically, in the task of chart generation, we parse and execute code derived from LLM responses and establish rigorous test cases to assess the accuracy of the generated charts, with a particular focus on the precision of y-axis fields, employing the pass@$1$ metric (Chen et al., 2021a) for evaluation.
4.4 Main Results
@@
Table 4 showcases the main results of over 30 advanced advanced LLMs on the TableBench. GPT-4 outperforms other models in numerous tasks, demonstrating superior performance across complex reasoning scenarios. Particularly in numerical computation and analytical tasks, GPT-4 maintains a commendable level of performance. TableLLM finetuned on the open-source models with TableInstruct achieves a performance level comparable to GPT-3.5, significantly validating the effectiveness of our training data. Despite these advancements, humans still surpass all LLMs in these tasks. Nevertheless, certain advanced LLMs, especially those employing proprietary approaches, demonstrate potential in these scenarios.
Category Analysis
Coarse-grained (Table 4) and fine-grained (Figure 5) experimental results reveal that most models perform commendably in fact-based reasoning tasks, indicating their proficiency in this area. However, challenges arise in numerical reasoning tasks due to the complexity of mathematical computations, especially complex calculations such as aggregation, which require multiple intermediate steps to reach the final answer. Data analysis tasks necessitate more intricate and comprehensive analytical skills, such as using correlation coefficients to analyze model relationships and employing linear regression functions to predict future trends, thereby imposing higher demands on the overall reasoning abilities of LLMs. The task of chart generation poses the greatest challenge, requiring significant coding skills and strict adherence to instructions. Notably, smaller-sized models exhibit significant deficiencies in chart generation tasks, highlighting their limitations in utilizing code to handle complex tasks.
Reasoning Methods Analysis
As illustrated in Table 4, those methods incorporating reasoning steps demonstrate a clear advantage on TableBench compared to methods that derive conclusions directly. The TCoT method exhibits stable and superior performance across various dimensions. The PoT method delivers commendable results in purely numerical computations, particularly in chart generation, but falls short in textual reasoning. We investigate the factors contributing to the suboptimal performance of the PoT method and find that the code execution success rate constrains the performance, as we only conduct a single generation and execute the code without employing any strategy for code correction. Even for the best-performing GPT4-Turbo, the executable code ratio is only 78.67%. This indicates that the PoT method requires LLMs with significant code-generation capabilities and instruction-following ability. However, it also underscores the substantial potential of the PoT method. Conversely, the SCoT method adapts effectively in scenarios requiring a combination of numerical and textual reasoning, such as analytical tasks, achieving a balanced yet modest overall performance. The performance of SCoT falls short of expectations due to its reliance on simulated outcomes rather than executing actual code, consequently amplifying the error margin.
@@
PCC with Auto Metric 1 0.981 0.995
@@
4.5 Consistency of Evaluation Methods
Despite constraints imposed on the output format and the standardization of ground truth annotations, the ROUGE-L metric may not fully capture the real performance due to the inherent flexibility in the outputs of LLMs. Both GPT-4 and human judgment are conducted, as shown in Table 5, to assess this potential bias. The Pearson Correlation Coefficient (Cohen et al., 2009) (PCC) is adopted to analyze the consistency across different evaluation methods. The results, as presented in the table, indicate a high level of agreement among these evaluating methods, demonstrating that the constraints are effective and our metric accurately reflects the real performance of LLMs on the TableBench.
5 Further Analysis
5.1 Instruct Following Analysis
@@
We observe that the performance trends of small-size LLMs across different reasoning methods differ from those observed in large-size models in Table 4. We find that the non-reasoning method (DP) performs better on small-size LLMs than reasoning-based methods. As shown in Figure 6, most models exhibit good instruction-following capabilities with the DP method due to the simpler instructions to follow. Conversely, small-size LLMs perform significantly worse with the PoT method, mainly due to their insufficient code generation capabilities, resulting in a lower rate of executable code generation. Additionally, the iterative symbolic reasoning steps required by the SCoT method pose considerable challenges for small-scale models.
@@
Figure 7: Data efficient of TableInstruct
5.2 Data Efficiency of TableInstruct
@@
In this section, we discuss the data efficiency of TableInstruct on the SFT process. We construct datasets of varying sizes by sampling from TableInstruct with sampling rates ranging from 0.2 to 0.6. Figure 7 visually depicts the relative performance at different sampling rates. Surprisingly, with only 60% of the samples, the model retains over 90% of the performance of the complete dataset. The full data provides the highest knowledge coverage, enabling the model to achieve optimal overall performance, comparable to GPT-3.5, with inference costs being only a fraction, indicating the high efficiency of TableInstruct.
6 Related Work
@@
Table QA (Mueller et al., 2019; Jin et al., 2022) has grown substantially, driven by the development of robust datasets that engage advanced algorithms in the tasks of semantic comprehension Huang et al. (2024); Yang et al. (2020, 2022b, 2022a); Li et al. (2023c, 2024); Bai et al. (2023); Yang et al. (2024b). These datasets function as significant milestones for enhancing table-centric semantic understanding. WTQ (Pasupat and Liang, 2015), SQA (Iyyer et al., 2017), and TabFact (Chen et al., 2019) set the cornerstone for Table QA research. They furnish benchmarks founded on question-answer pairs predicated on HTML tables sourced from Wikipedia. However, these datasets rely heavily on specific cell content from the table to formulate answers, which can not fully represent the multi-dimensional queries posed in real-world scenarios.
@@
Acknowledging this incongruity, some datasets have been introduced to bridge the gap. ToTTo (Parikh et al., 2020), OTTQA (Chen et al., 2020), and FeTaQA (Nan et al., 2022) step into the fore by providing free-form QA datasets. These datasets challenge models to generate answers that go beyond the table’s explicit content, thereby enhancing model performance to align with the free-form nature of real-world questions. FinQA (Chen et al., 2021b) and AIT-QA (Katsis et al., 2022) lay emphasis on numeric-focused queries. These datasets predominantly target financial tables, suggesting complex reasoning challenges that necessitate models to not only interpret but also to compute and extract nuanced information precisely. Further diversifying the landscape, datasets such as WikiSQL (Zhong et al., 2017), Spider (Yu et al., 2018), and Bird (Li et al., 2023a) introduce logical expressions as supervisory signals to train Table QA models, discreting reasoning capabilities through logic-based problem-solving. Despite the significant advancements made by LLMs in TableQA (Li et al., 2022; Singha et al., 2023; Li et al., 2023b; Lei et al., 2023), there is still a critical need for benchmarks that reflect the reasoning complexity encountered in real-world tabular data scenarios. TableBench, a comprehensive and complex benchmark, incorporates real-world complexities into its evaluation scenarios, effectively addressing the limitations of existing benchmarks
7 Conclusion
@@
We acknowledge the following limitations of this study: (1) This paper primarily focuses on the reasoning complexity of table questions. It does not extensively explore the inherent complexities of the tables themselves, such as large-size tables and hierarchical tabulation. (2) Tabular data in image formats, which are also prevalent in real-world applications, are not discussed in this paper.
Ethics Statement
This research adheres to ethical guidelines for AI development. We aim to enhance the capabilities of LLMs while acknowledging potential risks such as bias, misuse, and privacy concerns. To mitigate these, we advocate for transparency, rigorous bias testing, robust security measures, and human oversight in AI applications. Our goal is to contribute positively to the field and encourage responsible AI development and deployment.

### v2

@@ Introduction
we adopt Rouge-L (Lin 2004) to assess the quality of the generated answers by measuring the n-gram overlap with reference answers. In the PoT method, we enforce a specific format for the executable code outputs and evaluate the final answer with the ROUGE-L metric, ensuring alignment with other reasoning methodologies. Specifically, in the task of chart generation, we parse and execute code derived from LLM responses and establish rigorous test cases to assess the accuracy of the generated charts, with a particular focus on the precision of y-axis fields, employing the pass@$1$ metric (Chen et al. 2021a) for evaluation.
Fact Checking Num-Reasoning Data Analysis Visualization Overall
TCoT SCoT PoT TCoT SCoT PoT TCoT SCoT PoT TCoT SCoT PoT TCoT SCoT PoT
Human Performance 94.3 87.1 82.1 86.3 85.91
Open-source In Context Learning Methods
Llama2-7B 34.99 27.47 3.61 6.70 4.63 3.95 14.31 12.49 1.56 0.00 0.00 0.00 12.36 9.95 2.76
CodeLlama-7B 33.06 12.34 19.44 5.43 2.99 13.31 16.16 17.06 1.79 0.00 0.00 0.00 12.30 9.28 8.85
Gemma-7B 27.63 10.07 21.62 6.78 2.91 10.45 20.33 11.76 6.74 0.00 0.00 2.00 13.96 6.97 9.81
Mistral-7B 50.45 40.56 6.25 8.73 5.77 2.60 21.99 21.12 1.19 0.00 0.00 0.00 17.86 15.11 2.35
Deepseek-Coder-7B 22.92 27.48 48.98 6.45 5.61 34.66 18.73 20.72 18.17 8.00 18.00 18.00 13.10 14.58 28.89
CodeQwen1.5-7B 30.56 32.94 0.00 6.24 5.68 0.00 27.04 22.47 0.00 2.00 0.00 0.00 16.80 14.85 0.00
Qwen1.5-7B 56.08 53.53 39.23 11.30 10.99 20.40 24.77 22.96 7.66 0.00 0.00 0.00 20.70 19.65 16.29
Qwen2-7B 57.70 57.52 0.00 16.09 16.65 0.76 24.02 21.50 0.38 0.00 4.00 2.00 22.77 22.26 0.60
StructLM-7B 47.72 64.06 13.54 9.55 19.97 11.48 19.59 23.83 4.38 0.00 0.00 0.00 17.06 25.21 8.30
MAP-Neo-7B 32.70 33.22 0.00 7.23 6.46 0.00 21.85 14.38 0.44 0.00 0.00 4.00 15.26 12.03 0.40
Llama3-8B 38.32 72.53 13.94 22.02 17.33 19.50 30.15 30.75 9.31 0.00 0.00 10.00 25.71 27.59 14.43
Llama3.1-8B 47.89 36.29 30.38 11.26 13.77 17.24 15.78 14.82 8.86 8.00 0.00 8.00 16.76 15.81 14.88
Llama2-13B 48.47 32.69 3.03 15.83 6.79 4.48 22.04 17.16 3.19 0.00 0.00 0.00 20.86 13.25 3.61
StructLM-13B 26.28 64.49 1.04 12.30 17.38 0.00 20.70 18.41 0.28 0.00 0.00 0.00 16.35 21.94 0.21
WizardLM-13B 53.93 46.01 8.33 13.79 16.52 14.79 22.61 20.16 3.73 0.00 0.00 4.00 20.75 20.23 9.12
Qwen1.5-14B 40.83 61.92 44.38 10.29 15.01 28.20 22.99 29.24 10.33 2.00 8.00 2.00 18.03 25.14 21.48
Qwen1.5-32B 64.99 67.86 49.01 19.13 21.15 34.01 24.27 28.29 17.43 4.00 8.00 8.00 25.38 28.30 27.79
Deepseek-Coder-33B 48.27 54.34 33.12 9.41 12.69 32.60 9.09 21.70 19.97 0.00 0.00 24.00 13.01 19.92 27.20
CodeLlama-34B 64.39 58.28 5.90 13.10 13.30 4.20 19.23 15.28 0.53 0.00 0.00 2.00 20.24 18.19 2.88
StructLM-34B 19.10 30.21 27.74 15.36 9.03 14.45 20.74 17.92 5.38 0.00 0.00 2.00 16.93 14.37 11.61
Mixtral-8x7B 54.54 56.01 35.86 16.80 16.05 26.23 24.69 25.67 13.96 2.00 0.00 6.00 23.14 23.24 21.32
Qwen1.5-72B 71.27 67.03 33.16 19.01 16.68 20.85 26.63 27.33 13.03 2.00 8.00 14.00 26.66 25.80 18.65
Qwen2-72B 72.50 71.13 56.37 36.97 31.81 41.33 32.20 31.85 22.36 20.00 14.00 12.00 38.13 35.14 33.91
Qwen1.5-110B 74.87 69.80 53.55 29.81 23.33 36.83 27.34 29.32 18.38 14.29 12.00 24.00 32.81 30.10 30.77
Llama3-70B 73.88 75.44 60.64 37.64 28.87 36.59 37.47 34.06 26.11 4.00 6.00 10.00 39.59 34.48 33.59
Llama3.1-70B 76.32 77.65 59.05 44.89 38.93 34.04 35.88 33.87 23.15 26.00 6.00 34.00 43.85 39.22 32.52
Close-source In Context Learning Methods
GPT-3.5-Turbo 59.95 75.68 60.92 23.45 23.16 42.09 34.40 32.54 30.25 10.00 4.00 38.00 30.85 31.41 39.34
Qwen-Max 70.48 68.21 50.42 32.83 25.06 32.80 27.87 30.98 19.41 18.00 8.00 30.00 34.26 31.04 29.39
Yi-Large 71.41 66.08 13.19 40.18 23.20 15.25 29.22 22.59 5.97 26.00 26.00 6.00 38.57 27.82 10.90
GLM-4 67.93 73.59 31.49 34.01 26.18 25.46 30.47 31.54 25.34 8.00 14.00 22.00 34.80 32.76 25.92
Deepseek-Chat-V2 72.41 69.89 57.48 50.07 38.96 45.96 38.07 34.44 30.37 40.00 24.00 46.00 47.22 39.63 41.07
Deepseek-Coder-V2 73.90 70.17 63.00 47.22 38.24 47.26 33.09 31.82 31.56 40.00 26.00 44.00 44.26 38.57 42.75
GPT-4-Turbo 75.92 77.62 70.08 53.01 44.31 49.31 41.03 36.52 34.63 62.00 32.00 48.00 51.32 44.26 45.69
GPT-4o 72.63 71.01 62.31 54.46 42.20 47.83 38.90 34.65 30.03 56.00 38.00 54.00 50.39 42.22 42.92
Open-Source Fine-Tuning Methods
[…]
Table 3: The main results of advanced LLMs on TableBench are presented alongside human performance. All methods involving code generation and computation, particularly in the chart generation task, execute code only once to derive the final answer. The overall results represent a weighted average of performance across different categories.
Main Results
@@
Table 3 showcases the main results of over 30 advanced advanced LLMs on the TableBench. GPT-4 outperforms other models in numerous tasks, demonstrating superior performance across complex reasoning scenarios. Particularly in numerical computation and analytical tasks, GPT-4 maintains a commendable level of performance. TableLLM finetuned on the open-source models with TableInstruct achieves a performance level comparable to GPT-3.5, significantly validating the effectiveness of our training data. Despite these advancements, humans still surpass all LLMs in these tasks. Nevertheless, certain advanced LLMs, especially those employing proprietary approaches, demonstrate potential in these scenarios. However, complex reasoning environments on tabular data still remain challenges.
@@
PCC with Auto Metric 1.000 0.981 0.995
@@
Category Analysis
Experimental results in Table 3 reveal that most models perform commendably in fact-based reasoning tasks, indicating their proficiency in this area. However, challenges arise in numerical reasoning tasks due to the complexity of mathematical computations, especially complex calculations such as aggregation, which require multiple intermediate steps to reach the final answer. Data analysis tasks necessitate more intricate and comprehensive analytical skills, such as using correlation coefficients to analyze model relationships and employing linear regression functions to predict future trends, thereby imposing higher demands on the overall reasoning abilities of LLMs. The task of chart generation poses the greatest challenge, requiring significant coding skills and strict adherence to instructions. Notably, smaller-sized models exhibit significant deficiencies in chart generation tasks, highlighting their limitations in utilizing code to handle complex tasks.
Reasoning Methods Analysis
As illustrated in Table 3, those methods incorporating reasoning steps demonstrate a clear advantage on TableBench compared to methods that derive conclusions directly. The TCoT method exhibits stable and superior performance across various dimensions. The PoT method delivers commendable results in purely numerical computations, particularly in chart generation, but falls short in textual reasoning. We investigate the factors contributing to the suboptimal performance of the PoT method and find that the code execution success rate constrains the performance, as we only conduct a single generation and execute the code without employing any strategy for code correction. Even for the best-performing GPT4-Turbo, the executable code ratio is only 78.67%. This indicates that the PoT method requires LLMs with significant code-generation capabilities and instruction-following ability. However, it also underscores the substantial potential of the PoT method. Conversely, the SCoT method adapts effectively in scenarios requiring a combination of numerical and textual reasoning, such as analytical tasks, achieving a balanced yet modest overall performance. The performance of SCoT falls short of expectations due to its reliance on simulated outcomes rather than executing actual code.
Consistency of Evaluation Methods
Despite constraints imposed on the output format and the standardization of ground truth annotations, the ROUGE-L metric may not fully capture the real performance due to the inherent flexibility in the outputs of LLMs. Both GPT-4 and human judgment are conducted, as shown in Table 4, to assess this potential bias. The Pearson Correlation Coefficient (Cohen et al. 2009) (PCC) is adopted to analyze the consistency across different evaluation methods. The results, as presented in the table, indicate a high level of agreement among these evaluating methods, demonstrating that the constraints are effective and our metric accurately reflects the real performance of LLMs on the TableBench.
Further Analysis
@@
Instruct Following Analysis
We observe that the performance trends of small-size LLMs across different reasoning methods differ from those observed in large-size models in Table 3. In further analysis, we introduce a comparison with non-reasoning methods, specifically focusing on Direct Prompting (DP), which provides solutions directly without intermediate reasoning steps. We find that the non-reasoning method (DP) performs better on small-size LLMs than reasoning-based methods. As shown in Figure 5, most models exhibit good instruction-following capabilities with the DP method due to the simpler instructions to follow. Conversely, small-size LLMs perform significantly worse with the PoT method, mainly due to their insufficient code generation capabilities, resulting in a lower rate of executable code generation. Additionally, the iterative symbolic reasoning steps required by the SCoT method pose considerable challenges for small-scale models.
@@
[…]
Data Efficiency of TableInstruct
@@
In this section, we discuss the data efficiency of TableInstruct on the SFT process. We construct datasets of varying sizes by sampling from TableInstruct with sampling rates ranging from 0.2 to 0.6. Figure 6 visually depicts the relative performance at different sampling rates. Surprisingly, with only 60% of the samples, the model retains over 90% of the performance of the complete dataset. We observe that the Llama-3-8B model requires fewer than 4,000 samples to surpass the performance of Qwen1.5-70B on the dataset, demonstrating that the TableInstruct corpus significantly enhances tabular reasoning in smaller models. The full data provides the highest knowledge coverage, enabling the model to achieve optimal overall performance, comparable to GPT-3.5, with inference costs being only a fraction, indicating the high efficiency of TableInstruct.
Related Work
@@
Table QA (Mueller et al. 2019; Jin et al. 2022) has grown substantially, driven by the development of robust datasets that engage advanced algorithms in the tasks of semantic comprehension (Huang et al. 2024; Li et al. 2023c, 2024b; Bai et al. 2023; Yang et al. 2024; Li et al. 2022a, 2024c, 2024a). These datasets function as significant milestones for enhancing table-centric semantic understanding. WTQ (Pasupat and Liang 2015), SQA (Iyyer, Yih, and Chang 2017), and TabFact (Chen et al. 2020) set the cornerstone for Table QA research. They furnish benchmarks founded on question-answer pairs predicated on HTML tables sourced from Wikipedia. However, these datasets rely heavily on specific cell content from the table to formulate answers, which can not fully represent the multi-dimensional queries posed in real-world scenarios.
@@
Acknowledging this incongruity, some datasets have been introduced to bridge the gap. ToTTo (Parikh et al. 2020), OTTQA (Chen et al. 2021b), and FeTaQA (Nan et al. 2022) step into the fore by providing free-form QA datasets. These datasets challenge models to generate answers that go beyond the table’s explicit content, thereby enhancing model performance to align with the free-form nature of real-world questions. FinQA (Chen et al. 2021c) and AIT-QA (Katsis et al. 2022) lay emphasis on numeric-focused queries. These datasets predominantly target financial tables, suggesting complex reasoning challenges that necessitate models to not only interpret but also to compute and extract nuanced information precisely. Further diversifying the landscape, datasets such as WikiSQL (Zhong, Xiong, and Socher 2017), Spider (Yu et al. 2018), and Bird (Li et al. 2023a) introduce logical expressions as supervisory signals to train Table QA models, discreting reasoning capabilities through logic-based problem-solving. Despite the significant advancements made by LLMs in TableQA (Li et al. 2022b; Singha et al. 2023; Li et al. 2023b; Lei et al. 2023; He et al. 2023), there is still a critical need for benchmarks that reflect the reasoning complexity encountered in real-world tabular data scenarios. TableBench, a comprehensive and complex benchmark, incorporates real-world complexities into its evaluation scenarios, effectively addressing the limitations of existing benchmarks
Conclusion
@@
We acknowledge the following limitations of this study: (1) This paper mainly focuses on the reasoning complexity of table questions, which does not extensively explore the inherent complexities of the tables themselves. (2) Tabular data in image formats, which are also prevalent in real-world applications, are not discussed in this paper.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant Nos. 62276017, 62406033, U1636211, 61672081), and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLCCSE-2024ZX-18).